# Simplified Mamba with Disentangled Dependency Encoding for Long-Term Time Series Forecasting

### v2

@@ Abstract
Recent advances in deep learning have led to the development of numerous models for Long-term Time Series Forecasting (LTSF). However, most approaches still struggle to comprehensively capture reliable and informative dependencies inherent in time series data. In this paper, we identify and formally define three critical dependencies essential for improving forecasting accuracy: the order dependency and semantic dependency in the time dimension as well as cross-variate dependency in the variate dimension. Despite their significance, these dependencies are rarely considered holistically in existing models. Moreover, improper handling of these dependencies can introduce harmful noise that significantly impairs forecasting performance. To address these challenges, we explore the potential of Mamba for LTSF, highlighting its three key advantages to capture three dependencies, respectively. We further empirically observe that nonlinear activation functions used in vanilla Mamba are redundant for semantically sparse time series data. Therefore, we propose SAMBA, a Simplified Mamba with disentangled dependency encoding. Specifically, we first eliminate the nonlinearity of vanilla Mamba to make it more suitable for LTSF. Along this line, we propose a disentangled dependency encoding strategy to endow Mamba with efficient cross-variate dependency modeling capability while minimizing the interference between time and variate dimensions. We also provide rigorous theory as a justification for our design. Extensive experiments on nine real-world datasets demonstrate the effectiveness of SAMBA over state-of-the-art forecasting models.
1 Introduction
Long-term Time Series Forecasting (LTSF) plays a critical role in various real-world applications, such as early disaster warning and long-term energy scheduling (Zhou et al. (2021); Wu et al. (2021)). Accurate LTST often relies on extracting reliable dependencies from longer historical time horizons. For example, periodic and trend patterns are frequently leveraged in existing literature to guide model design and interpret results (Wu et al. (2021)). However, the success of simple linear model (Zeng et al. (2023)) and patching strategy in Transformer (Nie et al. (2022)) suggests that only capturing periodic and trend patterns may be insufficient for accurate forecasting. In this paper, we explore the predictability of time series by decomposing dependencies into three key aspects: the order dependency and semantic dependency in time dimension, as well as the cross-variate dependency. Specifically, order dependency represents the sequential nature among time points, semantic dependency captures the high-level temporal variation beyond surface numerical values, and cross-variate dependency refers to the inter-variate relationships in multi-variate time series data. This decomposition offers a new perspective for model design in LTSF.
In fact, most existing LTSF models fail to comprehensively consider both order and semantic dependencies. For example, linear models (e.g., DLinear (Zeng et al. (2023))) rely on point-wise mappings to preserve order dependency but are too simplistic to capture complex semantic dependency. In contrast, Transformer-based models are effective in learning semantic dependency through the self-attention mechanism (Vaswani et al. (2017)) and show improved performance when local temporal semantics is enriched using patching strategies(Nie et al. (2022)). However, they struggle with perceiving temporal order due to the permutation-invariant nature of self-attention, even with positional encodings (Zeng et al. (2023)). Therefore, capturing both order and semantic dependencies simultaneously remains an underexplored problem in LTSF.
Besides the order and semantic dependencies, cross-variate dependency is also crucial in many real-world scenarios, such as traffic flow, where significant correlations among variates have been observed (Han et al. (2024); Liu et al. (2020)). However, existing approaches that utilize cross-variate dependency (Channel Dependent, CD) frequently underperform compared to methods that treat each variate independently (Channel-Independent, CI) (Nie et al. (2022)). This phenomenon arises because previous CD approaches often encounter issues with over-smoothing and difficulty in fitting individual variate series (Chen et al. (2024)). Although a recent method proposes to perform CI modeling first before CD modeling to alleviate the issues (Liu et al. (2023)), our findings, from both empirical and theoretical, indicate that the less relevant temporal dependencies extracted from other variates can still harm the informativeness of the embedding produced for the target variate sequence. Consequently, effectively leveraging cross-variate dependency while maintaining the unique temporal dynamics and characteristics of each variate continues to be a challenging problem.
Recently, Mamba (Gu & Dao (2023)) has emerged as a strong contender to the transformer architecture in the field of natural language processing (NLP) (Park et al. (2024)). Given the similar sequential structure of time series and text data, Mamba also demonstrates great potential in solving LTST task. Through extensive experiments, we identify that Mamba presents three major advantages over Transformer for LTSF: (i) Mamba uses State Space Models (SSMs) to process sequences recursively, inherently capturing order dependency(Gu et al. (2021b)). (ii) Its selection mechanism, similar to self-attention, enables it to focus on or ignore particular inputs, making it well-suited for learning semantic dependency. (iii) Mamba’s near-linear complexity provides an efficiency advantage in encoding cross-variate dependency, which are typically computationally expensive when dealing with large amounts of variates.
Despite these advantages, directly adapting Mamba for LTSF is still non-trivial due to two key challenges: (i) Mamba was initially designed for NLP tasks, where semantic dependency is of primary importance. In contrast, LTSF typically emphasizes order dependency due to the sparsity of information in time series as opposed to natural language (Zeng et al. (2023)). (ii) Although Mamba can effectively capture both order and semantic dependencies, it lacks mechanisms for modeling cross-variate dependency, which is crucial for multivariate forecasting. Simply incorporating cross-variate dependency into Mamba using a CD strategy may exacerbate the issue by introducing irrelevant correlations, undermining the model’s effectiveness to handle individual variate series properly.
To overcome these challenges, we propose SAMBA, a Simplified Mamba with disentangled dependency encoding specifically tailored for LTSF. First, we remove the nonlinearities in vanilla Mamba, which we find to be redundant and prone to overfitting in the context of semantically sparse time series data, thereby improving its generalization ability. Furthermore, we introduce a theoretically sound disentangled encoding strategy that explicitly separates cross-time and cross-variate dependencies to avoid mutual interference between the two dimensions. We conduct extensive experiments on real-world forecasting benchmarks, demonstrating the superiority of SAMBA over the state-of-the-art methods. In particular, SAMBA proves universally effective across datasets with varying degrees of cross-variate dependency. Our contributions are summarized as follows.
@@ 1 Introduction
  • We identify and formally define three critical dependencies in time series data to guide the design of the LTSF models.
  • Our comprehensive analysis reveals two insights: (i) Compared to Linear model and Transformers, Mamba can effectively capture both order and semantic dependencies. (ii) Directly applying MLP, Transformer, and Mamba to LTSF can lead to overfitting issues due to the presence of nonlinearities. To mitigate this, we propose a simplified Mamba by removing nonlinearities and incorporating a theoretically sound disentangled encoding strategy that appropriately integrates cross-variate dependency to the model, enhancing the model’s global representation and predictive capabilities.
  • We empirically demonstrate that SAMBA achieves state-of-the-art performance across mainstream benchmarks. Our analysis also highlights the general utility of disentangled encoding, offering a universal strategy for developing future LTSF models.
2 Related works
@@
Deep learning has made significant advances in the field of natural language processing (Devlin et al. (2018)) and speech recognition (Dong et al. (2018)), inspiring researchers to repurpose these models for time series forecasting (Zhou et al. (2021); Wu et al. (2021)). The mainstream methods for LTSF currently include linear models and Transformer-based models. In particular, because of the strong modeling capability of the transformer, numerous works have utilized transformers to achieve superior performance. However, applying Transformers to long sequence forecasting presents two major challenges: 1) computational overhead due to quadratic space-time complexity and 2) the difficulty of capturing long-term dependencies. Early approaches concentrated on modifying the architectural design to reduce complexity. For example, LogTrans (Li et al. (2019)) introduced convolutional self-attention to capture local information while utilizing LogSparse techniques to reduce complexity. Informer (Zhou et al. (2021)) leveraged a self-attention distillation mechanism to extend the input length and proposed ProbSparse self-attention to optimize computation efficiency. Furthermore, Autoformer (Wu et al. (2021)) selected a decomposed structural framework to disintegrate time and replaced traditional self-attention mechanisms with autocorrelation mechanisms to reduce complexity.
@@
Recently, linear models have demonstrated superior performance with fewer parameters and higher efficiency, prompting researchers to question whether transformers are suitable for long sequence forecasting (Zeng et al. (2023)). Researchers have begun exploring how to let transformers exhibit their powerful performance in scenarios like natural language processing and speech based on the inherent properties of time series. For example, considering the distribution shift problem inherent in time series, methods like Non-stationary transformers (Liu et al. (2022b)) have applied techniques to stabilize and reduce the inherent distribution shift of time series. Given the insufficiency of semantic information at individual time points, the PatchTST (Nie et al. (2022)) model, by patching to extract local information, addresses the issue of insufficient short-time series semantic content, thus enriching the semantic content of each token.
@@
Recent research shows that utilizing the CI strategy to achieve promising results (Nie et al. (2022)), inspiring researchers to explore a new encoding method that can consider both intra-variate and inter-variate interactions (Chen et al. (2023)). Crossformer (Zhang & Yan (2022)) adopted a cross-variate encoding method and a two-stage attention mechanism that considers both time and variate relationships. iTransformer (Liu et al. (2023)) applied a data inversion method, considering the time series corresponding to each variate as a token and then capturing the relationships between variates using self-attention mechanisms. CARD (Wang et al. (2024)) explicitly models both cross-time dependency and cross-variate dependency, where the latter directly mixes these dependencies at each model layer. As a result, CARD still suffers from the issue mentioned in the introduction, which is the inappropriate mixing of cross-time and cross-variate dependencies.
Unlike previous work, we explore a new architecture more suitable for LTSF—Mamba (Gu & Dao (2023)), with modifications to better fit the LTSF scenarios. Moreover, different from previous approaches to serially encoding time and variate relationships, we propose a parallel encoding method that decouples cross-time and cross-variate dependencies. This disentanglement ensures that the encoding process of each dependency does not adversely affect each other, leading to state-of-the-art (SOTA) performance on datasets with varying degrees of variate relationships.
3 Problem Formulation
@@
The LTSF aims to learn a mapping function $\mathcal{F}(\cdot)$ that forecasts the temporal evolution of $N$ variates in the future $S$ time steps based on the observations in the past $T$ time steps, i.e.,
$\mathcal{F}:\mathbb{R}^{N\times T}\rightarrow\mathbb{R}^{N\times S},\ \mathbf{X}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{T})\mapsto\mathbf{Y}=(\mathbf{x}_{T+1},\ldots,\mathbf{x}_{T+S}),$ (1)
@@
$I(c_{t};c_{s}\mid t,s)>I(c_{t};c_{s}),$ (2)
where $I(\cdot;\cdot)$ ($I(\cdot;\cdot\mid\cdot)$) denotes the (conditional) mutual information (Cover (1999)). For example, the increasing trend of daily temperatures in the past week helps predict that the trend will likely continue in the future.
@@
$H(\bm{c}_{T+1:T+S}|\mathcal{O}(\bm{c}_{1:T}),\mathcal{S}(\bm{c}_{1:T}))<H(\bm{c}_{T+1:T+S}|\mathcal{O}(\bm{c}_{1:T})).$ (3)
@@
4 Empirical Exploration of Mamba for LTSF
An effective LTSF model should be able to take advantage of the three types of dependencies defined in Section 3. We first conduct in-depth ablation experiments to evaluate three models: two classical LTSF models, Linear model and Transformer, and the third, Mamba, in capturing these dependencies. The details of the implementation of the following experiments are provided in Appendix B.4.
4.1 Why Mamba? A Deep Dive into Its Suitability for LTSF
To validate Mamba’s suitablity for LTSF, we evaluate its capability to capture both order dependency and semantic dependency. The evaluation of Mamba’s performance in capturing cross-variate dependency is provided in Appendix C. To assess the models’ ability to capture order dependency, we first make the following assumption.
Assumption 1. The performance of an effective order dependency learner is significantly influenced by the order of the input sequence (Zeng et al. (2023)).
According to Assumption 1, a greater performance variation indicates a greater reliance on order dependency. Our results in Table 1 show that both the Linear model and Mamba surpass the Transformer in terms of capturing order dependency. This is because the Linear model and Mamba that use the SSM-based approach are permutation-variant, i.e., regarding the order. Their sensitivity to sequence order is stronger than that of the Transformer, which relies on the permutation-invariant self-attention mechanism. This increased sensitivity accounts for the greater performance decrease observed for the linear model and Mamba, as illustrated in Table 1.
Table 1: Order dependency analysis on ETTm1 dataset (Zhou et al. (2021)). Based on Assumption 1, we compare each model’s performance variation before and after randomly shuffling the temporal order of testing data points. O.MSE and O.MAE are evaluated in the original test set. S.MSE and S.MAE are evaluated in the shuffling test set.
  Models Linear Model Mamba Transformer
Metrics
O.MSE
S.MSE
O.MAE
S.MAE
O.MSE
S.MSE
O.MAE
S.MAE
O.MSE
S.MSE
O.MAE
S.MAE
96 0.383 0.988 0.400 0.697 0.517 0.922 0.508 0.688 0.643 0.884 0.575 0.643
192 0.413 0.986 0.415 0.697 0.575 0.931 0.546 0.699 0.805 1.01 0.664 0.730
336 0.441 0.987 0.435 0.698 0.730 0.957 0.634 0.703 0.882 1.12 0.737 0.817
720 0.497 0.992 0.469 0.704 0.873 0.973 0.704 0.723 0.928 1.12 0.752 0.80
Avg. Drop - 127.97% - 62.55% - 40.37% - 17.60% - 22.40% - 6.55%

### v3

# SDE: A Simplified and Disentangled Dependency Encoding Framework for State Space Models in Time Series Forecasting

@@ Abstract.
In recent years, advancements in deep learning have spurred the development of numerous models for Long-term Time Series Forecasting (LTSF). However, most existing approaches struggle to fully capture the complex and structured dependencies inherent in time series data. In this work, we identify and formally define three critical dependencies that are fundamental to forecasting accuracy: order dependency and semantic dependency along the temporal dimension, as well as cross-variate dependency across the feature dimension. These dependencies are often treated in isolation, and improper handling can introduce noise and degrade forecasting performance. To bridge this gap, we investigate the potential of State Space Models (SSMs) for LTSF and emphasize their inherent advantages in capturing these essential dependencies. Additionally, we empirically observe that excessive nonlinearity in conventional SSMs introduce redundancy when applied to semantically sparse time series data. Motivated by this insight, we propose SDE (Simplified and Disentangled Dependency Encoding), a novel framework designed to enhance the capability of SSMs for LTSF. Specifically, we first eliminate unnecessary nonlinearities in vanilla SSMs, thereby improving the suitability for time series forecasting. Building on this foundation, we introduce a disentangled encoding strategy, which empowers SSMs to efficiently model cross-variate dependencies while mitigating interference between the temporal and feature dimensions. Furthermore, we provide rigorous theoretical justifications to substantiate our design choices. Extensive experiments on nine real-world benchmark datasets demonstrate that SDE-enhanced SSMs consistently outperform state-of-the-art time series forecasting models.
Time series forecasting; Disentangled encoding; State space model
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737119isbn: 979-8-4007-1454-2/2025/08ccs: Information systems Spatial-temporal systemsccs: Computing methodologies Artificial intelligence
KDD Availability Link:
The source code of this paper has been made publicly available at https://github.com/YukinoAsuna/SAMBA.
1. Introduction
Long-term Time Series Forecasting (LTSF) plays a crucial role in numerous real-world applications, such as early disaster warning systems and long-term energy planning (Zhou et al., 2021; Wu et al., 2021; Yeh et al., 2024; Wang et al., 2024a). The effectiveness of LTSF models relies heavily on their ability to capture meaningful dependencies from extensive historical observations. As a result, traditional approaches often leverage periodicity and trend patterns to guide model design and enhance interpretability (Wu et al., 2021). However, the recent success of linear models (Zeng et al., 2023) and patching strategies in Transformers (Nie et al., 2022) suggests that existing formulations may overlook deeper and more essential factors that drive predictability in time series.
In this work, we revisit the foundations of time series modeling by decomposing temporal dependencies into three key aspects: (i) order dependency, which reflects the sequential structure across time steps; (ii) semantic dependency, which captures high-level temporal patterns beyond surface numerical values; and (iii) cross-variate dependency, which models interactions among variables in multivariate time series. We argue that effective LTSF requires careful modeling of all three dependencies. This provides a unified lens for analyzing existing models and their limitations.
Currently, deep LTSF models typically fall into three paradigms: (i) linear- and MLP-based models (Li et al., 2023; Zeng et al., 2023; Das et al., 2023; Chen et al., 2023), (ii) Transformer-based models (Wu et al., 2021; Zhou et al., 2021; Liu et al., 2023; Nie et al., 2022; Liu et al., 2024), and (iii) State Space Models (SSMs) (e.g. RNN-based models) (Lin et al., 2023; Jia et al., 2024; Jhin et al., 2024). The majority of existing LTSF approaches rely on either MLP/linear models or Transformer-based architectures. However, both of them exhibit fundamental limitations in capturing order and semantic dependencies. Linear models (e.g., DLinear (Zeng et al., 2023)) excel at preserving temporal order through point-wise mappings, but struggle to capture semantic-level patterns due to the restricted capacity. Transformer models (Vaswani et al., 2017), in contrast, are effective in learning semantic dependency via attention mechanisms, especially when enhanced with patching strategies (Nie et al., 2022), but often lack the structural inductive bias needed to preserve sequential order (Zeng et al., 2023; Tan et al., 2024). Despite the use of positional encodings, the self-attention mechanism remains permutation-invariant, limiting its temporal sensitivity (Tan et al., 2024).
On the other hand, SSMs offer a promising alternative. Their recursive structure naturally preserves order dependencies (Gu et al., 2021), and recent advances in deep SSMs have significantly improved their capacity to capture semantic patterns. Nevertheless, conventional SSM-based implementations, particularly through Recurrent Neural Networks (RNNs), still suffer from gradient vanishing and limited parallelism when handling long-range sequential dependencies, making them less competitive for LTSF tasks (Jia et al., 2024; Lin et al., 2023).
To address these issues, two main research directions emerged: (i) improving the efficiency and stability of RNN-style architectures (Lin et al., 2023; Jia et al., 2024), and (ii) developing new SSM variants tailored for long-range dependencies, such as the recently proposed Mamba (Gu and Dao, 2023; Park et al., 2024). While these methods show promise in various sequence tasks, the direct application to LTSF faces two non-trivial challenges: (1) Order-semantic dependency rebalance: Mamba and similar SSMs are primarily designed for language modeling, where semantic richness dominates. In contrast, time series are often low in information density (Zeng et al., 2023) and require stronger modeling of order dependency. Naively applying such architectures may lead to imbalance, overfitting to semantic noise while ignoring essential sequential patterns. (2) Cross-variate dependency modeling: Multivariate forecasting relies on capturing variable interactions. Most SSMs lack explicit mechanisms for modeling cross-variate dependencies (Han et al., 2024; Liu et al., 2020). However, existing cross-variate modeling strategies (e.g., Channel-Dependent (CD) methods) often underperform compared to Channel-Independent (CI) approaches (Liu et al., 2023). This discrepancy arises from the over-smoothing effect in CD models, losing the unique temporal features of individual variables (Chen et al., 2024). While hybrid designs (e.g., CI-then-CD) have been proposed (Liu et al., 2023), our empirical and theoretical analysis demonstrate that they still introduce redundant information from weakly correlated variables, ultimately compromising forecasting performance.
To tackle the above challenges, we propose SDE, a Simplified and Disentangled Dependency Encoding framework to enhance SSMs for LTSF. Specifically, SDE consists of two major components: simplification and disentangled encoding strategy. (1) Simplification mechanism: We eliminate unnecessary nonlinear components from existing SSMs, such as RNNs and Mamba. Our empirical and theoretical analysis show that such nonlinearities often cause overfitting in time series with low semantic complexity. By simplifying these components, we improve generalization and robustness across datasets. (2) Disentangled encoding strategy: We introduce a theoretically grounded disentangled encoding strategy that explicitly separates temporal and cross-variate dependencies. This disentangled design reduces mutual interference between dimensions and is universally applicable across both SSM and Transformer backbones, enabling flexible integration. Through extensive experiments on multiple real-world datasets, we demonstrate that our method—SDE-Mamba, a Mamba variant enhanced by our framework—consistently achieves state-of-the-art performance. Notably, SDE-Mamba adapts well to datasets with varying levels of cross-variate correlation, underscoring its robustness and generality. Our major contributions are summarized as follows:
@@ 1. Introduction
  • We systematically identify and formally define three critical types of dependencies in time series data, i.e., order dependency, semantic dependency, and cross-variate dependency, to guide the design of LTSF models.
  • We derive two key insights: (i) SSMs, compared to linear models and Transformers, exhibit superior capability in capturing both order and semantic dependencies; (ii) vanilla SSMs suffer from overfitting when applied to LTSF due to excessive nonlinear operations.
  • In response, we propose the SDE framework, which consists of two core components: (i) a simplification mechanism applicable to both RNN- and Mamba-based SSMs, removing excessive nonlinear operations to improve generalization; and (ii) a theoretically grounded disentangled encoding strategy that effectively integrates cross-variate dependencies while preserving the dynamics of individual variates.
  • Extensive experiments validate the effectiveness of our approach, with SDE-Mamba achieving SOTA performance across multiple LTSF benchmarks. Furthermore, our analysis showcase the broad applicability of disentangled encoding strategy, offering a generalizable framework for future LTSF model development.
2. Related Works
@@
Deep learning has made significant advances in the field of natural language processing (Devlin et al., 2018), intelligent Internet of Things (Xiao et al., 2023), speech recognition (Dong et al., 2018), inspiring researchers to repurpose these models for time series forecasting (Zhou et al., 2021); (Wu et al., 2021). The mainstream methods for LTSF currently include linear models and Transformer-based models. In particular, because of the strong modeling capability of the transformer, numerous works have utilized transformers to achieve superior performance. However, applying Transformers to long sequence forecasting presents two major challenges: 1) computational overhead due to quadratic space-time complexity and 2) the difficulty of capturing long-term dependencies. Early approaches concentrated on modifying the architectural design to reduce complexity. For example, LogTrans (Li et al., 2019) introduced convolutional self-attention to capture local information while utilizing LogSparse techniques to reduce complexity. Informer (Zhou et al., 2021) leveraged a self-attention distillation mechanism to extend the input length and proposed ProbSparse self-attention to optimize computation efficiency. Furthermore, Autoformer (Wu et al., 2021) selected a decomposed structural framework to disintegrate time and replaced traditional self-attention mechanisms with autocorrelation mechanisms to reduce complexity.
@@
Recently, linear models have demonstrated superior performance with fewer parameters and higher efficiency, prompting researchers to question whether transformers are suitable for long sequence forecasting (Zeng et al., 2023). Researchers have begun exploring how to let transformers exhibit their powerful performance in scenarios like natural language processing and speech based on the inherent properties of time series. For example, considering the distribution shift problem inherent in time series, methods like Non-stationary transformers (Liu et al., 2022a) have applied techniques to stabilize and reduce the inherent distribution shift of time series. Given the insufficiency of semantic information at individual time points, the PatchTST (Nie et al., 2022) model, by patching to extract local information, addresses the issue of insufficient short-time series semantic content, thus enriching the semantic content of each token.
@@
Recent research shows that utilizing the CI strategy to achieve promising results (Nie et al., 2022), inspiring researchers to explore a new encoding method that can consider both intra-variate and inter-variate interactions (Chen et al., 2023). Crossformer (Zhang and Yan, 2022) adopted a cross-variate encoding method and a two-stage attention mechanism that considers both time and variate relationships. iTransformer (Liu et al., 2023) applied a data inversion method, considering the time series corresponding to each variate as a token and then capturing the relationships between variates using self-attention mechanisms. CARD (Wang et al., 2024b) explicitly models both cross-time dependency and cross-variate dependency, where the latter directly mixes these dependencies at each model layer. As a result, CARD still suffers from the issue mentioned in the introduction, which is the inappropriate mixing of cross-time and cross-variate dependencies.
Unlike previous studies, we conduct a comprehensive analysis of the advantages of state space models (SSMs) in LTSF and propose SDE, an enhanced framework that fully leverages their strengths. In contrast to existing approaches, SDE-Mamba explicitly states-of-the-art (SOTA) performance across datasets with varying degrees of intervariate dependencies.
3. Problem Formulation
Problem Statement 1.
@@ 3. Problem Formulation
Long-term Time Series Forecasting (LTSF). The goal is to learn a mapping function $\mathcal{F}(\cdot)$ that forecasts the temporal evolution of $N$ variates in the future $S$ time steps based on the observations in the past $T$ time steps, i.e.,
(1) $\mathcal{F}:\mathbb{R}^{N\times T}\to\mathbb{R}^{N\times S},\quad\mathbf{X}\mapsto\mathbf{Y},$
(2) $\mathbf{X}=(\mathbf{x}_{1},\dots,\mathbf{x}_{T}),\quad\mathbf{Y}=(\mathbf{x}_{T+1},\dots,\mathbf{x}_{T+S}).$
@@
(3) $I(c_{t};c_{s}\mid t,s)>I(c_{t};c_{s}),$
where $I(\cdot;\cdot)$ ($I(\cdot;\cdot\mid\cdot)$) denotes the (conditional) mutual information (Cover, 1999). For example, the increasing trend of daily temperatures in the past week helps predict that the trend will likely continue.
@@
(4) $H(\bm{c}_{T+1:T+S}|\mathcal{O}(\bm{c}_{1:T}),\mathcal{S}(\bm{c}_{1:T}))<H(\bm{c}_{T+1:T+S}|\mathcal{O}(\bm{c}_{1:T})).$
@@
4. Empirical Exploration of State Space Model for LTSF
An effective LTSF model should be able to take advantage of the three types of dependencies defined in Section 3. We first conduct in-depth ablation experiments to evaluate four models: two classical LTSF models, Linear and Transformer, as well as two state space models in capturing these dependencies.
4.1. Why State Space model? A Deep Dive into Its Suitability for LTSF
To verify the applicability of SSMs to LTSF, we selected SegRNN (Lin et al., 2023) and Mamba (Gu and Dao, 2023) as two representative SSMs to evaluate SSM’s ability to capture sequential dependencies and semantic dependencies. To assess the models’ ability to capture order dependency, we first make the following assumption.
Assumption 1. The performance of an effective order dependency learner is significantly influenced by the order of the input sequence (Zeng et al., 2023).
According to Assumption 1, a greater performance variation indicates a greater reliance on order dependency. In addition, we also consider the SOTA model, iTransformer as one of our baslines. It is worth noting that iTransformer encodes temporal relationships using a Linear Model and then uses a Transformer encoder to capture cross-variate dependencies. Ideally, it is expected to be sensitive to order as well. Our results in Table 3 show that Linear model, Mamba and SegRNN surpass the Transformer in terms of capturing order dependency. This is because the Linear model, as well as the state-space models Mamba and SegRNN, which process sequences recursively, are all permutation-variant, i.e., sensitive to the order. Their sensitivity to sequence order is stronger than that of the Transformer, which relies on the permutation-invariant self-attention mechanism. This increased sensitivity accounts for the greater performance decrease observed for the linear model and Mamba, as illustrated in Table 3.